# Prompting in Practice: Investigating Software Developers’ Use of Generative AI Tools

### v1

@@ 3.3.3. Issue Frequency
Our final reliability analysis examined how frequently developers encountered specific issues (given to them) when using GenAI tools across different SE contexts. Participants rated twelve task-specific issues on a five-point scale from “Always” (1) to “Never” (5), where lower mean scores indicate more frequent issue occurrence. Results are visualized in Figure 7.
Debugging emerges as the most problematic SE task, with suggested bug fixes missing the root cause having the highest issue frequency (2.34, n=32). This issue received four “Always” responses (the highest count for any absolute rating across all issues, despite only appearing for those who selected the third-most popular task), indicating that GenAI’s diagnostic capabilities remain a critical hurdle. Refactoring and code generation follow in severity, with unintended changes to behavior (2.52, n=23) and correctness/compatibility issues in code (2.58, n=66) representing significant challenges. Documentation and code review showed the lowest reported issue frequencies, with developers noticing inaccurate or incomplete technical documentation (3.06, n=34) and false positives and hallucinations (3.00, n=23) least often. This aligns with our previous findings about the lower complexity of these tasks.
Finding 14: Respondents reported experiencing the most issues when using GenAI for debugging, particularly when it would miss the root cause of the problem. Code generation and refactoring were also noted pain points.
@@ 4.1. Developing Proficiency with Generative AI
Our results show a positive relationship between a developer’s perceived GenAI proficiency and the number of tasks they complete with GenAI (Table 1), suggesting that task diversity and GenAI proficiency may mutually reinforce each other. This could potentially reflect how confidence encourages developers to explore GenAI usage in new contexts, or could also demonstrate that using GenAI on a wide range of tasks increases proficiency.
@@
4.2. Limitations and Practical Implications
@@
While these results provide some direction for the critical problems GenAI engineers should focus on solving, others point toward alternative ways to improve reliability and user experience. Table 6 shows how the number of GenAI exchanges increases with a user’s skill, and users tend to prefer Iterative conversations to Comprehensive ones (Section 3.2.4). Accordingly, high-profile tools usually facilitate iteration in some respect, most obviously by allowing users to send multiple prompts throughout a conversation. However, there are other ways to facilitate iteration, such as allowing users to edit earlier prompts to create conversation branches.
Our findings indicate that daily use makes productivity improvements significantly more likely (Section 3.1.2), and also correlates with GenAI use on a broader range of tasks. This suggests that the best way for individual developers to realize the benefits of these new tools is to carefully consider which tasks in their workflow could benefit from utilizing GenAI, then purposefully learn about and apply GenAI there. Gradual adjustment may be necessary to ensure consistent performance and build an understanding of how best to navigate AI limitations.
@@ 4.3. High Technique Awareness, Low Adoption
Our survey population had an overwhelming familiarity with various prompting techniques, with their awareness of the Meta Language Creation technique being the lowest (74%). However, despite an overall high awareness of prompting techniques, there was a lackluster adoption. For example, even though 92% (66/72) of respondents indicated familiarity with the Output Style technique, only 82% of those (54) actually used it at all, and of those only 54% (29) used it regularly. The adoption percentages are worse for the remaining techniques, and regular usages are well below 50%. This is particularly interesting given that awareness of research outcomes is typically one of the biggest hurdles for tool and technique adoption, meaning that the issue must lie elsewhere. It may be that either developers do not find these techniques helpful or that they are unaware of how to effectively use them. The results in Table 4 suggest that decades of work experience may assist veteran developers in finding use cases for prompting techniques, since this was the group most likely to adopt these techniques. However, flexibility may also be a crucial component to effective use, as potentially demonstrated by the intermediate developers who had the highest frequency of regular technique usage. Accordingly, future research can further explore the pain points that developers encounter when applying these techniques to their workflows. Additionally, a comprehensive taxonomy and/or usage guide could potentially assist developers in more effectively using these prompting techniques or understanding how to apply them to their particular use cases.
@@ 5. Threats to Validity
Internal Validity. To mitigate bias, we employed an iterative open-coding methodology for analyzing open-ended survey responses. We also followed best practices when formulating our survey questions in order to avoid confusing and/or biasing language, and conducted a small pilot study with graduate students to test the content, clarity, and length of our questionnaire. We recruited participants from a variety of sources (e.g., our professional networks, university organizations, etc.) to ensure we received a broad range of perspectives and experiences, but we are aware of the threat of self-selection bias.
External Validity. The conclusions of this study only apply to those participants that completed our survey and cannot be overly generalized to the larger developer population. However, given the themes and trends that emerged, we believe that many developers will share similar experiences to those reported. Ultimately, our goal was never to claim generalizability, but instead to explore the emerging landscape of how developers are interacting with GenAI tools for SE tasks.
Construct Validity. This work relies on self-reported developers’ perceptions, which can be subjective and influenced by recall and social desirability biases. We sought to mitigate this threat by corroborating and anchoring subjective assessments such as proficiency with more concrete metrics captured within the survey, including the breadth of AI-assisted tasks participants perform and their frequency of tool use, and avoiding making definite claims about practices. The demographic information provided in our study was supplied by respondents and cannot be independently verified.
@@ 6. Related Work
While many distinct techniques have been researched and compiled, there has been no investigation comparing developers’ preferred prompting strategies and interaction methods across multiple categories of software engineering tasks, leaving information about how these tools are used in practice relatively scarce. Our study addresses this gap with a comprehensive analysis of prompting and conversation practices in six distinct SE domains.
@@ 7. Conclusion
We conducted a comprehensive survey of 72 software developers to investigate how they integrate GenAI tools into software engineering workflows. Our findings reveal primary adoption of GenAI for code generation, while use for related tasks such as documentation and debugging strongly correlates with a developer’s AI proficiency and usage frequency. Users tend to prefer iterative conversations to achieve their goals, with consistent single-shot prompting remaining an impossibility. Significant reliability challenges remain for complex coding tasks, particularly debugging.

### v2

@@ 3.3.3. Issue Frequency
Our final reliability analysis examined how frequently practitioners encountered specific issues (given to them) when using GenAI tools across different SE contexts. Participants rated twelve task-specific issues on a five-point scale from “Always” (1) to “Never” (5), where lower mean scores indicate more frequent issue occurrence. Results are visualized in Figure 7.
Debugging emerges as the most problematic SE task, with suggested bug fixes missing the root cause having the highest issue frequency (2.34, n=32). This issue received four “Always” responses (the highest count for any absolute rating across all issues, despite only appearing for those who selected the task), indicating that GenAI’s diagnostic capabilities remain a critical hurdle. Refactoring and code generation follow in severity, with unintended changes to behavior (2.52, n=23) and correctness/compatibility issues in code (2.58, n=66) representing significant challenges. Documentation and code review showed the lowest reported issue frequencies, with practitioners noticing inaccurate or incomplete technical documentation (3.06, n=34) and false positives and hallucinations (3.00, n=23) least often. This aligns with our previous findings about the lower complexity of these tasks.
Finding 13: Respondents reported experiencing the most issues when using GenAI for debugging, particularly when it fails to provide root causes. Code generation and refactoring were also noted pain points.
@@ 4.1. Developing Proficiency with Generative AI
Our results show a positive relationship between a practitioner’s perceived GenAI proficiency and the number of tasks they complete with GenAI (Table 1), suggesting that task diversity and GenAI proficiency may mutually reinforce each other. This could potentially reflect how confidence encourages practitioners to explore GenAI usage in new contexts, or could also demonstrate that using GenAI on a wide range of tasks increases proficiency.
@@
4.2. GenAI Limitations & Practical Implications
@@
While these results provide some direction for the critical problems GenAI engineers should focus on solving, others point toward alternative ways to improve reliability and user experience. Table 6 shows how the number of GenAI exchanges increases with a user’s skill, and users tend to prefer iterative conversations (Section 3.2.4). Accordingly, high-profile tools usually facilitate iteration in some respect, most obviously by allowing users to send multiple prompts throughout a conversation. However, there are other ways to facilitate iteration, such as allowing users to edit earlier prompts to create conversation branches.
Our findings indicate that daily use makes productivity improvements significantly more likely (Section 3.1.2), and also correlates with GenAI use on a broader range of tasks. This suggests that the best way for individual practitioners to realize the benefits of these new tools is to carefully consider which tasks in their workflow could benefit from utilizing GenAI, then purposefully learn about and apply GenAI there. Gradual adjustment may be necessary to ensure consistent performance and build an understanding of how best to navigate AI limitations.
@@ 4.3. High Technique Awareness, Low Adoption
Our survey population had an overwhelming familiarity with various prompting techniques, with their awareness of the Meta Language Creation technique being the lowest (74%). However, despite an overall high awareness, there was a lackluster adoption. For example, even though 92% (66/72) of respondents indicated familiarity with the Output Style technique, only 82% of those (54) actually used it at all, and of those only 54% (29) used it regularly. The adoption percentages are worse for the remaining techniques, and regular usages are well below 50%. This is particularly interesting given that awareness of research outcomes is typically one of the biggest hurdles for tool and technique adoption, meaning that the issue must lie elsewhere. It may be that either practitioners do not find these techniques helpful or that they are unaware of how to effectively use them. The results in Table 4 suggest that decades of work experience may assist veteran practitioners in finding use cases for prompting techniques, since this was the group most likely to adopt these techniques. However, flexibility may also be a crucial component to effective use, as potentially demonstrated by the intermediate practitioners who had the highest frequency of regular technique usage. Accordingly, future research can further explore the pain points that practitioners encounter when applying these techniques to their workflows. Additionally, a comprehensive taxonomy and/or usage guide could potentially assist practitioners in more effectively using these prompting techniques or understanding how to apply them to their particular use cases.
@@ 5. Threats to Validity
Internal Validity. To mitigate possible coding biases, we used a rigorous iterative open-coding methodology to analyze open-ended survey responses. We also followed the best practices when formulating our survey questions in order to avoid confusing and/or biasing language, and conducted a small pilot study with graduate students to test the clarity and length of our questionnaire. We recruited participants from a variety of sources (our professional networks, university organizations, etc.) to ensure we received a broad range of perspectives and experiences, but we are aware of the threat of self-selection bias.
External Validity. Study conclusions apply only to participants who completed our survey and cannot be generalized to larger practitioner populations. However, given the themes and trends that emerged, we believe that many practitioners will share similar experiences to those reported. Ultimately, our goal was never to claim generalizability, but to explore the emerging landscape of how practitioners are interacting with GenAI tools for SE tasks.
Construct Validity. This work relies on self-reported practitioner perceptions, which can be subjective and influenced by recall and social desirability biases. We sought to mitigate this threat by corroborating and anchoring subjective assessments such as proficiency with more concrete metrics captured within the survey, including the breadth of AI-assisted tasks participants perform and their frequency of tool use, and avoiding making definite claims about practices.
@@ 6. Related Work
While many distinct techniques have been researched and compiled, there has been no investigation comparing practitioners’ preferred prompting strategies and interaction methods across multiple categories of software engineering tasks, leaving information about how these tools are used in practice relatively scarce. Our study addresses this gap with a comprehensive analysis of prompting and conversation practices in six distinct SE domains.
@@ 7. Conclusion
We conducted a comprehensive survey of 72 software practitioners to investigate how they integrate GenAI tools into software engineering workflows. Our findings reveal primary adoption of GenAI for code generation, while use for related tasks such as documentation and debugging strongly correlates with a practitioner’s AI proficiency and usage frequency. Users tend to prefer iterative conversations to achieve their goals, with consistent single-shot prompting remaining an impossibility. Significant reliability challenges remain for complex coding tasks, particularly debugging.